# Confidence as Forecast: A Decision-Theoretic Interpretation of Confidence Intervals

### v1

@@ 1.1 Background
The ”it either covers, or it doesn’t” statement has been the source of confusion, consternation, and frustration for beginning statistics students for a long time, likely since it was first made, and has led to a number of not-quite-satisfying claims about how to interpret CIs in the applied literature. Introductory papers and applied practitioners occasionally claim that constructed intervals retain their nominal coverage probability [9, 8], critiques in response claim that they retain none [12, 4, 7, 17], and still others seem so befuddled by it all that they suggest what intervals are trying to estimate are the interval endpoints themselves (rather than $\theta$) [14]. More often than not, the ensuing debates–which can be rather spirited [13, 18, 11]–end with trenches being drawn along philosophical lines, with frequentists defending claims that are sometimes, on their face, rather awkward (e.g., retreating to a von Mises-style infinite hypothetical frequentism without acknowledging that his program, taken seriously, undermines the foundational pillars of Kolmogorov-style mathematical statistics [5]), and Bayesians wondering why the frequentists will not just switch to Bayesianism, put priors over $\theta$, and construct credible intervals instead.
In the sections below, I hope to offer a principled frequentist reading of the concept of ”confidence” and its associated intervals that resolves some of this confusion. More precisely, the notion I build toward recasts the notion of coverage probability as having three layers, rather than a single one: the first being the event-level degenerate conditional in $\{0,1\}$ determining whether an interval covers; the second being the design-level coverage guarantee of $1-\alpha$ that averages those conditionals over the randomness in $X$; and the third, or the notion of ”confidence”, being a predictive probability, or model-based forecast, of empirical coverage with whatever information the statistician may have at hand. Under this view, we have clear bounds on what we can say about a particular interval ex post (e.g., we might predict that it covers $\theta$ with probability $1-\alpha$), and, in some cases, we also have mathematical justification for updating our forecast in light of new evidence, for example, when we see the ”trivial” interval $\infty$$\infty$], where we would very sensibly switch our prediction to $1$, since coverage is certain. Separating our probabilistic forecasts from the design-level coverage guarantee generally washes away criticisms of CIs as being uninterpretable (with respect to the coverage event and not the actual value of $\theta$), and, as I show below, it generally respects the frequentist machinery Neyman used to build his theory (even if he might have preferred not to see it that way).
@@ 1.2 Paper overview
The remainder of the paper is organized as follows. Section 2 presents a though experiment showing that design-level coverage probability can in fact be used to guide decision-making based on constructed intervals ex post. Section 3 introduces the relevant standard formalisms for frequentist statistical inference, and it formalizes the notion of ”confidence” as a predictive probability, or model-based estimate, of interval coverage that applies both ex ante and ex post. Section 4 shows how this predictive notion might apply to CI-based inference by revisiting a commonly-cited thought experiment designed to disprove the notion that confidence has any meaning ex post, and Section 5 concludes with a recap of the main results from earlier sections; a discussion of pedagogical strategies for teaching both CI theory and its applications; and some suggested potential directions for future research.
@@ 2 A thought experiment: ”Monty’s Hell”
“Take a look at these 3 plastic cups. Under each one, I have placed a handwritten note indicating a range of dollar amounts—for example, \$10 to \$20. On a separate sheet of paper that you can’t see, I’ve written down a single dollar amount that falls under one, and only one, of the ranges hidden by the cups. I will shuffle the cups for a while and then let you pick one to overturn. I’ll then remove one of the cups you didn’t pick from the table and give you the opportunity to switch your original choice. After you make your final choice, if the range on the paper beneath your cup contains the hidden amount on my other sheet, I’ll pay you the hidden amount; if not, you’ll pay me half of the hidden amount. The game costs \$5 to play, and the hidden amount is at least \$10—are you in?”
@@ 2.1 The parallels with CIs
This thought experiment was designed as a direct analog to CIs, and there are two main parallels to note. First is that, as with CIs capturing $\theta$, each cup winning is decided solely by whether its interval captures the hidden dollar amount. Second, we have that the winning value, like $\theta$, is a fixed-but-unknown constant, with no probability distribution being placed over its value. These two facts fix whether a cup wins or loses, but, as with CIs, we have no way of knowing which. The uncomfortable dissonance here is that from both Bayesian and frequentist simulation-based analyses of the original Monty Hall problem under its standard setup, we know that switching is, indeed, the optimal strategy. Does Neyman’s interpretation also lead to success? The answer in this case is an emphatic ”no”: if we follow either of the options he gives us for interpreting CIs after choosing our first cup, we are guaranteed to lose money. The degenerate conditional in $\{0,1\}$ would freeze us in our tracks, since we now have no basis for weighing whether to switch and thus decide to stay; and stating ”this interval covers” would also do the same, since we now have a clear reason not to switch, as we are treating our first cup as the one that wins. In both cases, the long run would keep the chances of winning at $1/3$, and we would lose an average of $\mathdollar 5$ per game, rather than taking advantage of the $2/3$ chances offered by the switch. In other words, the two Neyman-style moves—refusing to assign a nondegenerate probability to coverage, or simply declaring that our original cup “covers”—are not only philosophically awkward; rather, they are always outperformed by the strategy that treats the design-level success probability as a forecast.
@@ 3.1 What the model gives us
$E_{\theta}[Z(X)]=1-\alpha,\qquad\text{that is,}\qquad Z(X)\sim\mathrm{Bernoulli}(1-\alpha).$ (4)
@@
The two layers are linked by the usual tower property. Combining (4) and (5),
$E_{\theta}[Z(X)]\;=\;E_{\theta}\!\Big[E_{\theta}\big(Z(X)\,\big|\,X\big)\Big]\;=\;E_{\theta}\big[Z(X)\big]\;=\;1-\alpha.$ (6)
Thus the nondegenerate coverage probability $1-\alpha$ lives at the design level, as the unconditional mean of the coverage indicator, while the degenerate $\{0,1\}$ values arise upon conditioning on the data $X$. Both are internal to the same model: they are simply $Z$ viewed with respect to different $\sigma$–fields. In particular, they apply to all intervals generated by the procedure, regardless of whether we have already observed the corresponding data or not.
@@ 3.3.2 Minimizing risk post-trial
$\mathbb{E}_{\theta}\!\left[S\bigl(q(X),Z_{\theta}\bigr)\,\big|\,\mathcal{G}\right].$
@@
$q^{*}(X)\;=\;\mathbb{E}_{\theta}\!\left[Z_{\theta}\,\middle|\,\mathcal{G}\right]\;=\;P_{\theta}\bigl(\theta\in I(X)\,\big|\,\mathcal{G}\bigr).$ (8)
@@
$\mathbb{E}_{\theta}\!\left[Z_{\theta}\,\middle|\,\mathcal{G}\right]\;=\;1-\alpha\quad\text{almost surely},$ (9)
@@
$P_{\theta}\bigl(\theta\in I(X)\mid\mathcal{G}\bigr)\neq 1-\alpha,$
@@
recovering the oracle’s degenerate view discussed in Section 3.1. This “forecast” is perfect for an oracle that knows whether the interval covers, but a non‑omniscient statistician who insists on the slogan “either it covers or it does not” is in a different position: they know only that the true coverage indicator takes values in $\{0,1\}$, but they still must name a single real number $q\in[0,1]$ when scored. Any operational rule $Q$ that implements this stance and only ever reports 0 or $1$, without access to $Z_{\theta}$ itself, has strictly larger expected loss than the design‑based forecast $q\equiv p=1-\alpha$ under any strictly proper Bernoulli scoring rule. Formally, if $Z_{\theta}\sim\mathrm{Bernoulli}(p)$ with $p=1-\alpha$ and $S$ is strictly proper, then for every $\{0,1\}$–valued forecast rule $Q$,
$E_{\theta}\bigl[S(Q,Z_{\theta})\bigr]>E_{\theta}\bigl[S(p,Z_{\theta})\bigr],\qquad 0<p<1,$
because $E_{\theta}[S(q,Z_{\theta})]$ is uniquely minimized at $q=p$ and $p\notin\{0,1\}$. In this sense, the vague “forecast” $\{0,1\}$ is strictly dominated, in the usual scoring‑rule sense, by the sharp design‑level forecast $1-\alpha$.
@@ 3.4.1 Unbounded location–scale models
In standard unbounded, translation-invariant models with pivot-based CIs, interval coverage depends on a pivot, while interval width is ancillary (see the Supplemental Methods to see how this works in the case of estimating a normal mean with a $t$ interval). In such designs, neither the absolute location of the interval nor any $\theta$-free function of its geometry carries additional design-based information about coverage, and the optimal pre- and post-trial forecast for $\mathbf{1}\{\theta\in I(X)\}$ under any strictly proper scoring rule is simply the nominal level $1-\alpha$. Many, if not most, real-world estimation problems will fall under this category.
@@ 4.2.2 Beating the constant forecasts with conditional coverage
The reason Morey et al. present this thought experiment, however, is naturally not to extol the virtues of the confidence level as a constant forecast: indeed, it is primarily to show that, given the underlying model, a constant design-level forecast leads one to make rather awkward statements about coverage ex post. For example, if we use the NP procedure to construct an interval for a sample and notice that, after construction, the interval only spans $25\%$ of the $10$-meter length of the hull, should we still say it covered the hatch with $50\%$ probability? In a sense, yes, that is exactly true–as shown in Section 3, the model gives us two layers of probability that apply to each and every interval, one of them being $1-\alpha$–although it does sound very strange to say, and it requires us to ignore the actual values (and relative width) of the realized interval, which throws away some information.
@@
Unsurprisingly, the conditional forecasts beat the constant design-level forecast in both cases, with the Brier score dropping from $0.25$ to about $0.11$ for the NP interval and to about $0.18$ for the UMP interval. The scores are instructive in another way as well. In this setup, the two procedures share the same coverage indicator $Z_{\theta}$: by definition, whenever the NP interval covers $\theta$, the UMP interval also covers, and vice versa. The difference lies in the $\theta$-free statistic used for forecasting coverage. For the NP procedure, we condition on its relative width $D$; for the UMP procedure, we condition on its (shorter) relative width
@@ 4.2.3 The bit about nesting
To see such reasoning in action, let us pick a new pair of procedures, one being the UMP procedure from above, and the other being what Morey et al. describe as the ”sampling distribution” (SD) procedure, defined as $\bar{x}\pm(5-5/\sqrt{2})$. Under simulation, these have similar marginal coverage probabilities, at $49,998\,/\,1e5\approx 0.500$ for the former, and $50,094\,/\,1e5\approx 0.501$ for the latter. The SD procedure does not overlap with the UMP procedure in the same way as the NP procedure, though, and the joint coverage probability under simulation turns out to be a tad higher at $58,605\,/1e5\,\approx 0.586$. From the start, then, we know there should be no reason to consider the composite procedure as having the same chance of capturing $\theta$ as each procedure individually–the design-level coverage probability nearly $9\%$ higher, presumably because we are casting a wider net on average than with either procedure alone.
What, then, about nesting? Can we use it to our advantage for forecasting? As it turns out, coverage probability does change substantially when the intervals are nested: when the SD intervals are nested inside the UMP intervals, the probability that either of them covers goes up to $0.793$, and when the roles are reversed, it drops down below the design-level joint coverage to $0.441$ (note that, because the intervals share the same midpoint by design, one is always nested inside of the other–all that changes is which one is where). These probabilities are also stable across the simulation configurations, and so they appear to be $\theta$-free. To update our ex-ante forecast from the design-level joint value $p_{\mathrm{joint}}\approx 0.586$ to something different ex post, we now have two main choices: use the conditional coverage probabilities $P(Cover\,|\,\mathrm{SD}\subset\mathrm{UMP})$ and $P(Cover\,|\,\mathrm{UMP}\subset\mathrm{SD})$ given above, or use the finer-grained versions of these where use coverage probability conditioned on the relative width of whichever is the wider of the two intervals (i.e., the outer one). Table 2 shows the Brier scores for these forecasts, relative to the outcome that either of the intervals cover.
@@
Again, a forecast of always $1$ leads to the worst loss at $0.414$; moving to the constant design-level joint forecast $q\equiv p_{\mathrm{joint}}\approx 0.586$ lowers that by slightly less than half to $0.243$. Using conditional coverage probabilities under the two nesting conditions lowers a touch more to $0.213$, and combining that information with each procedure’s marginal conditional coverage probability (relative to itself, not to either interval covering) leads to the lowest score at $0.208$. These results offer some mild empirical evidence in favor of the general principle raised in Section 3 that conditioning on more information leads to a better forecast.
Before leaving the submarine example behind, we can return to the apparent logical paradox raised at the beginning of this subsection about the potentially-missing coverage probability between the inner and outer intervals in a nested pair. Across the simulations, the inner SD interval misses when the outer UMP interval covers with an average probability of $0.085$, and vice-versa (results were again stable across all combinations of $\theta$ and scale). This probability fills the ”gaps” between the inner and outer nested interval in both cases and are not, thankfully, 0, as Morey et al. had wondered whether they might be. It is also approximately the difference between either of the procedures’ marginal coverage probability of $0.500$ and their joint coverage probability of $0.586$, a sensible and welcome result.
@@ 5.1. The utility of confidence as a forecast
Treating ”confidence” as a probabilistic forecast gets us a few, hopefully helpful, things for free when using Neyman’s machinery for inference. As an interpretation for some number of realized intervals, it gives us a principled way to estimate the number of them that covered, for example, by using $1-\alpha$ as $p$ to calculate a binomial mean. This benefit most notably applies to the case of a single realized interval, where we can now happily say both that the interval either covers, or it does not (conditioned on its endpoints), and that it has some intermediate probability of covering the parameter (conditioned on the group coverage of other intervals like it). It also allows for a straightforward interpretation of groups of intervals that very plainly would not cover the parameter with probability $1-\alpha$, since we have the mathematical support for updating our forecast based on any coverage-relevant information they may contain. As touched on above, the latter serves as a response to the numerous counter-examples in the literature intended to show that ”confidence” has no coherent meaning ex post, and the former serves as formal support for Neyman’s original intention for CIs as having estimable coverage properties in conducting repeated (empirical) experiments. Although these two seem the most valuable, treating ”confidence” as a forecast also helps keep clear what we mean when we talk about ”coverage probability”, which can now be seen as both a design-level unconditional probability under the model and a information-relative conditional probability based on whatever we happen to observe when using a confidence procedure.

### v2

@@ 1.1 Background
The ”it either covers, or it doesn’t” statement has been the source of confusion, consternation, and frustration for beginning statistics students for a long time, perhaps since it was first made, and has led to a number of not-quite-satisfying claims about how to interpret CIs in the applied literature. Introductory papers and applied practitioners occasionally claim that constructed intervals retain their nominal coverage probability [9, 7], critiques in response claim that they retain none [12, 4, 6, 17], and still others suggest what intervals are trying to estimate are the interval endpoints themselves, rather than $\theta$ [14]. More often than not, the ensuing debates–which can be rather spirited [13, 18, 11]–end at philosophical impasses, with frequentists and Bayesians each appealing to foundational commitments that the other side does not share. .
In the sections below, I hope to offer a principled frequentist reading of the concept of ”confidence” and its associated intervals that aims to resolve some of this confusion (see also [8] for a broader conceptual argument about ex-post probability statements) . More precisely, I treat coverage probability as having three layers, rather than a single one: the first being the event-level degenerate conditional in $\{0,1\}$ determining whether an interval covers; the second being the design-level coverage guarantee of $1-\alpha$ that averages those conditionals over the randomness in $X$; and the third, or the notion of ”confidence”, being a predictive probability, or model-based forecast, of empirical coverage with whatever information the statistician may have at hand. Under this view, we have clear bounds on what we can say about a particular interval ex post (e.g., we might predict that it covers $\theta$ with probability $1-\alpha$), and, in some cases, we also have mathematical justification for updating our forecast in light of relevant features of the realized interval, for example, when we see the ”trivial” interval $\infty$$\infty$], where we would very sensibly switch our prediction to $1$, since coverage is certain. Separating our probabilistic forecasts from the design-level coverage guarantee helps dissolve some complaints of CIs as being uninterpretable (with respect to the coverage event and not the actual value of $\theta$), and, as I show below, it generally respects the frequentist machinery Neyman used to build his theory (even if he might not have intended it to be used quite that way).
@@ 1.2 Paper overview
The remainder of the paper is organized as follows. Section 2 presents a thought experiment showing that design-level coverage probability can in fact be used to guide decision-making based on constructed intervals ex post. Section 3 introduces the relevant standard formalisms for frequentist statistical inference, and it formalizes the notion of ”confidence” as a predictive probability, or model-based estimate, of interval coverage that applies both ex ante and ex post. Section 4 shows how this predictive notion might apply to CI-based inference by revisiting a commonly-cited thought experiment designed to disprove the notion that confidence has any meaning ex post, and Section 5 concludes with a recap of the main results from earlier sections; a discussion of pedagogical strategies for teaching both CI theory and its applications; and some suggested potential directions for future research.
@@ 2 A thought experiment: ”Monty’s Hell”
“Take a look at these 3 plastic cups. Under each one, I have placed a handwritten note indicating a range of dollar amounts—for example, \$10 to \$20. On a separate sheet of paper that you can’t see, I’ve written down a single dollar amount that falls under one, and only one, of the ranges hidden by the cups. I will shuffle the cups for a while and then let you pick one to overturn. I’ll then remove one of the cups you didn’t pick that I know does not contain the winning range; if I have to choose between two such cups, you’ll flip a fair coin (provided by me) to decide which one I remove. I’ll then give you the opportunity to switch your original choice. After you make your final choice, if the range on the paper beneath your cup contains the hidden amount on my other sheet, I’ll pay you the hidden amount; if not, you’ll pay me half of the hidden amount. The game costs \$5 to play, and the hidden amount is at least \$10—are you in?”
@@ 2.1 The parallels with CIs
This thought experiment was designed as a direct analog to CIs, and there are two main parallels to note. First is that, as with CIs capturing $\theta$, each cup winning is decided solely by whether its interval captures the hidden dollar amount. Second, we have that the winning value, like $\theta$, is a fixed-but-unknown constant, with no probability distribution being placed over its value. These two facts fix whether a cup wins or loses, but, as with CIs, we have no way of knowing which. The uncomfortable dissonance here is that from both Bayesian and frequentist simulation-based analyses of the original Monty Hall problem under its standard setup, we know that switching is, indeed, the optimal strategy. Does Neyman’s interpretation also lead to success? The answer in this case is ”no”: if we follow either of the options he gives us for interpreting CIs after choosing our first cup, we are guaranteed to lose money. The degenerate conditional in $\{0,1\}$ would freeze us in our tracks, since we now have no basis for weighing whether to switch and thus decide to stay; and stating ”this interval covers” would also do the same, since we now have a clear reason not to switch, as we are treating our first cup as the one that wins. In both cases, the long run would keep the chances of winning at $1/3$, and we would lose an average of $\mathdollar 5$ per game, rather than taking advantage of the $2/3$ chances offered by the switch. In other words, the two Neyman-style moves—refusing to assign a nondegenerate probability to coverage, or simply declaring that our original cup “covers”—are not only philosophically awkward; rather, they are strictly outperformed by the strategy that treats the design-level success probability as a forecast.
@@ 3.1 What the model gives us
$E_{\theta}[Z(X)]=1-\alpha,$ (4)
so that, under $P_{\theta}$, $Z(X)$ is Bernoulli with success probability $1-\alpha$.
@@
The two layers are linked by the usual tower property:
[…]
where the inner expectation is the degenerate conditional from (6). Thus, the design-level coverage probability $1-\alpha$ is recovered by averaging those $\{0,1\}$ values over the sampling distribution of $X$. Both layers are internal to the same model: they are simply $Z$ viewed with respect to different $\sigma$–fields. In particular, they apply to all intervals generated by the procedure, regardless of whether we have already observed the corresponding data or not.
@@ 3.2.2 Minimizing risk post-trial
$E_{\theta}\!\Big[S\!\big(q(X),\,Z_{\theta}\big)\,\Big|\,\mathcal{G}\Big].$
@@
$q^{*}(X)=E_{\theta}[Z_{\theta}\mid\mathcal{G}]=P_{\theta}\!\big(\theta\in I(X)\,\big|\,\mathcal{G}\big).$ (9)
@@
$E_{\theta}[Z_{\theta}\mid\mathcal{G}]=1-\alpha\quad\text{almost surely},$ (10)
@@
$P_{\theta}\!\big(\theta\in I(X)\mid\mathcal{G}\big)\neq 1-\alpha,$
@@
[…]
@@ 3.3.1 Unbounded location–scale models
In standard unbounded, translation-invariant models with pivot-based CIs, the coverage event is determined by a pivot, while interval width is a $\theta$-free feature of the realized interval. In such designs, neither the absolute location of the interval nor any $\theta$-free function of its geometry carries additional design-based information about coverage, and the optimal pre- and post-trial forecast for $\mathbf{1}\{\theta\in I(X)\}$ under any strictly proper scoring rule is simply the nominal level $1-\alpha$. Many familiar estimation problems will fall under this category.
@@ 4.2.2 Beating the constant forecasts with conditional coverage
The reason Morey et al. present this thought experiment, however, is naturally not to extol the virtues of the confidence level as a constant forecast: indeed, it is primarily to show that, given the underlying model, a constant design-level forecast leads one to make rather awkward statements about coverage ex post. For example, if we use the NP procedure to construct an interval for a sample and notice that, after construction, the interval only spans $25\%$ of the $10$-meter length of the hull, should we still say it covered the hatch with $50\%$ probability? In a sense, yes, that is exactly true–as shown in Section 3, the model gives us two layers of probability that apply to each and every interval, one of them being $1-\alpha$–although it may feel unsatisfying to say, and it requires us to ignore the actual values (and relative width) of the realized interval, which throws away some information.
@@
Unsurprisingly, the conditional forecasts beat the constant design-level forecast in both cases, with the Brier score dropping from $0.25$ to about $0.12$ for the NP interval and to about $0.17$ for the UMP interval. The scores are instructive in another way as well. In this setup, the two procedures share the same coverage indicator $Z_{\theta}$: by definition, whenever the NP interval covers $\theta$, the UMP interval also covers, and vice versa. The difference lies in the $\theta$-free statistic used for forecasting coverage. For the NP procedure, we condition on its relative width $D$; for the UMP procedure, we condition on its (shorter) relative width
@@ 4.2.3 The bit about nesting
To see such reasoning in action, let us pick a new pair of procedures, one being the UMP procedure from above, and the other being what Morey et al. describe as the ”sampling distribution” (SD) procedure, defined as $\bar{x}\pm(5-5/\sqrt{2})$. Under simulation, these have similar marginal coverage probabilities, at $\approx 0.500$ for the former, and $\approx 0.501$ for the latter. The SD procedure does not overlap with the UMP procedure in the same way as the NP procedure, though, and the joint coverage probability under simulation turns out to be a tad higher at $\approx 0.585$. From the start, then, we know there should be no reason to consider the composite procedure as having the same chance of capturing $\theta$ as each procedure individually–the design-level coverage probability roughly $8.5\%$ higher, presumably because we are casting a wider net on average than with either procedure alone.
What, then, about nesting? Can we use it to our advantage for forecasting? As it turns out, coverage probability does change substantially when the intervals are nested: when the SD intervals are nested inside the UMP intervals, the probability that either of them covers goes up to $0.790$, and when the roles are reversed, it drops down below the design-level joint coverage to $0.440$ (note that, because the intervals share the same midpoint by design, one is always nested inside of the other–all that changes is which one is where). These probabilities are also stable across the simulation configurations, and so they appear to be $\theta$-free. To update our ex-ante forecast from the design-level joint value $p_{\mathrm{joint}}\approx 0.585$ to something different ex post, we now have two main choices: use the conditional coverage probabilities $P(Cover\,|\,\mathrm{SD}\subset\mathrm{UMP})$ and $P(Cover\,|\,\mathrm{UMP}\subset\mathrm{SD})$ given above, or use the finer-grained versions of these where use coverage probability conditioned on the relative width of whichever is the wider of the two intervals (i.e., the outer one). Table 2 shows the Brier scores for these forecasts, relative to the outcome that either of the intervals cover.
@@
Again, a forecast of always $1$ leads to the worst loss at $0.415$; moving to the constant design-level joint forecast $q\equiv p_{\mathrm{joint}}\approx 0.585$ lowers that by slightly less than half to $0.243$. Using conditional coverage probabilities under the two nesting conditions lowers a touch more to $0.213$, and combining that information with each procedure’s marginal conditional coverage probability (relative to itself, not to either interval covering) leads to the lowest score at $0.212$. These results offer some mild empirical evidence in favor of the general principle raised in Section 3 that conditioning on more information leads to a better forecast.
Before leaving the submarine example behind, we can return to the apparent logical paradox raised at the beginning of this subsection about the potentially-missing coverage probability between the inner and outer intervals in a nested pair. Across the simulations, the inner SD interval misses when the outer UMP interval covers with an average probability of $0.085$, and vice-versa (results were again stable across all combinations of $\theta$ and scale). This probability fills the ”gaps” between the inner and outer nested interval in both cases and are not, thankfully, 0, as Morey et al. had wondered whether they might be. It is also approximately the difference between either of the procedures’ marginal coverage probability of $0.500$ and their joint coverage probability of $0.585$, a sensible and welcome result.
4.2.4 Code availability
Code for the lost submarine simulations is available on GitHub at https://github.com/scotthlee/confidence-as-forecast.
@@ 5.1. The utility of confidence as a forecast
Treating ”confidence” as a probabilistic forecast gets us a few, hopefully helpful, things for free when using Neyman’s machinery for inference. As an interpretation for some number of realized intervals, it gives us a principled way to estimate the number of them that covered, for example, by using $1-\alpha$ as $p$ to calculate a binomial mean. This benefit most notably applies to the case of a single realized interval, where we can now happily say both that the interval either covers, or it does not (conditioned on its endpoints), and that it has some intermediate probability of covering the parameter (conditioned on the group coverage of other intervals like it). It also allows for a straightforward interpretation of groups of intervals that very plainly would not cover the parameter with probability $1-\alpha$, since we have the mathematical support for updating our forecast based on any coverage-relevant information they may contain. As touched on above, the latter serves as a response to the numerous counter-examples in the literature intended to show that ”confidence” has no coherent meaning ex post, and the former serves as formal support for Neyman’s original intention for CIs as having estimable coverage properties in conducting repeated (empirical) experiments. Although these two seem the most valuable, treating ”confidence” as a forecast also helps keep clear what we mean when we talk about ”coverage probability”, which can now be seen as both a design-level unconditional probability under the model and an information-relative conditional probability based on whatever we happen to observe when using a confidence procedure.